OpenStudy (anonymous):

sqrt of 4y^3 - sqrt of y^3

OpenStudy (anonymous):

\[\sqrt{4y ^{3}}-\sqrt{y ^{3}}=2\sqrt{y ^{3}}-\sqrt{y ^{3}}=\sqrt{y ^{3}}=y ^{3/2}\]

OpenStudy (anonymous):

Thats not one of the possible answers....

OpenStudy (anonymous):

\[y \ge0\]

OpenStudy (anonymous):